# A randomized small-block Lanczos method for large-scale null space computations

### v1

@@ 2.3 Principal angles and null space residual
To study the impact of perturbations on subspaces, we recall the definition and basic properties of principal angles. For later purposes, it will be helpful to admit subspaces of unequal dimension, as in [10, Section 6.4.3] and [38].
@@
Note that the principal angles in Definition 2.8 are sorted in increasing order, and it is often the largest principal angle $\angle_{k}(V_{0},V_{1})$ that is of interest. In particular, it is well known that $\sin\angle_{k}(V_{0},V_{1})=\lVert V_{0}V_{0}^{\mathsf{T}}-V_{1}V_{1}^{\mathsf{T}}\rVert$ holds when $k_{0}=k_{1}$, implying that the orthogonal projectors are close for small $\angle_{k}(V_{0},V_{1})$. The other principal angles give partial information; knowing that some of them are small allows us to conclude that $\mathcal{V}_{0}$ and $\mathcal{V}_{1}$ nearly share a common subspace. In the context of null spaces, the following result establishes a residual bound in terms of principal angles.
@@
Let $W_{t}$ be the right singular vectors corresponding to the largest $t$ singular values of $V_{0}^{\mathsf{T}}V_{1}$. Then
$\|(I-V_{0}V_{0}^{\mathsf{T}})V_{1}W_{t}\|^{2}$ $=$ $\max_{\|x\|=1}\|(I-V_{0}V_{0}^{\mathsf{T}})V_{1}W_{t}x\|^{2}$
$=$ $\max_{\|x\|=1}\big{(}\|V_{1}W_{t}x\|^{2}-\|V_{0}V_{0}^{\mathsf{T}}V_{1}W_{t}x\|^{2}\big{)}$
$=$ $1-\min_{\|x\|=1}\|V_{0}^{\mathsf{T}}V_{1}W_{t}x\|^{2}=1-\sigma_{t}^{2}=\sin^{2}\angle_{t}(V_{0},V_{1}).$
@@
We conclude this section with a useful characterization of the tangents of principal angles from [38, Theorem 3.1].
@@
Let $[V_{0},V_{\perp}]$ be an orthonormal matrix, where $V_{0}\in\mathbb{R}^{n\times k_{0}}$. Suppose $\Omega_{1}\in\mathbb{R}^{n\times k_{1}}$ has column full rank. If $\mathrm{rank}(V_{0}^{\mathsf{T}}\Omega_{1})=\mathrm{rank}(\Omega_{1})\leq k_{0}$, then
@@ 3.1 Convergence of single-vector Krylov subspaces for small eigenvalues
In the following, we analyze how well the Krylov subspace generated in line 3 of Algorithm 1 contains approximations to (portions of) the invariant subspace associated with the $N$ smallest eigenvalues of $B$. The analysis mimics analogous results obtained in [19] for low-rank approximation. Compared to classical results [28, Chapter VI], our analysis differs by considering the convergence to a whole subspace and leveraging the randomness of the initial vector. In the following theorem, the main result of this section; $\mathsf{Cheb}_{k}(\lambda)$ denotes the Chebyshev polynomial of degree $k$ defined as
@@
For a symmetric matrix $B\in\mathbb{R}^{n\times n}$ with eigenvalues $\lambda_{1}<\lambda_{2}<\dotsb<\lambda_{N}<\lambda_{N+1}\leq\dotsb\leq\lambda_{n}$, let $V_{N}\in\mathbb{R}^{n\times N}$ be an orthonormal basis of the invariant subspace associated with $\lambda_{1},\dotsc,\lambda_{N}$. Choose integers $t,\ell$ such that $1\leq t\leq N$ and $t\leq\ell\leq n$. Let $K_{\ell}=[\omega,B\omega,\dotsc,B^{\ell-1}\omega]$, where $\omega\in\mathbb{R}^{n}$ is a Gaussian random vector. Then, for any $10<\delta<$ 1, the inequality
@@
Following [19, Theorem 3], the ability of a (single-vector) Krylov subspace to capture a subspace is closely related to properties of Vandermonde matrices.
@@
where $\mathcal{P}_{t-1}$ denotes the vector space containing all polynomials of degree at most $t-1$. Choosing the Lagrange basis
@@
Because $V_{\perp}^{\mathsf{T}}\omega$ and $V_{N}^{\mathsf{T}}\omega$ are independent Gaussian random vectors, we can use concentration and anti-concentration results for chi-squared random variables from [15, Lemma 1] and [19, Lemma 4], respectively, to conclude that
@@
Our proof crucially relies on a simple but important observation from [19, (1.3)], which establishes the following equivalence between the single-vector Krylov subspace $\mathrm{span}\{K_{\ell}\}$ and a block Krylov subspace generated from a specific starting matrix:
(16) $\mathrm{span}\{K_{\ell}\}=\mathrm{span}\{\omega,B\omega,\dotsc,B^{\ell-1}\omega\}=\mathrm{span}\{K_{t},BK_{t},\dotsc,B^{\ell-t}K_{t}\},$
@@
and an analogous formula holds when $V_{N}$ is replaced by $V_{\perp}$. By assumption, $p(\Lambda_{N})$ is invertible and thus $V_{N}^{\mathsf{T}}KH_{K}$ has rank $t$. Additionally, using that $\mathrm{span}\{KH_{K}\}\subset\mathrm{span}\{K\}$ and the expression for $\tan\angle_{t}$ from Proposition 2.11, we obtain that
@@
$=\lVert p(\Lambda_{\perp})V_{\perp}^{\mathsf{T}}\Omega(V_{N}^{\mathsf{T}}\Omega)^{\dagger}p(\Lambda_{N})^{-1}\rVert$
$\leq\frac{\max\limits_{N+1\leq j\leq n}\lvert p(\lambda_{j})\rvert}{\min\limits_{1\leq i\leq N}\lvert p(\lambda_{i})\rvert}\lVert V_{\perp}^{\mathsf{T}}\Omega(V_{N}^{\mathsf{T}}\Omega)^{\dagger}\rVert=\frac{\max\limits_{N+1\leq j\leq n}\lvert p(\lambda_{j})\rvert}{\min\limits_{1\leq i\leq N}\lvert p(\lambda_{i})\rvert}\tan\angle_{t}(V_{N},\Omega).$
The proof is completed by setting $p(\lambda)=\mathsf{Cheb}_{\ell-t}\bigl{(}1+\frac{2(\lambda_{N+1}-\lambda)}{\lambda_{n}-\lambda_{N+1}}\bigr{)}$ and using that a Chebyshev polynomial is bounded by one in magnitude inside $[-1,1]$ and monotone outside that interval.
3.2 Convergence of Algorithm 1
@@
Under the setting of Theorem 2.6, let $V$ be the approximate null space basis returned after $\ell\geq N$ iterations of Algorithm 1. Let $\tilde{\lambda}_{t}^{(0)}$ be the $t$th smallest eigenvalue of $Z_{\ell}^{\mathsf{T}}A^{\mathsf{T}}AZ_{\ell}$, where $1\leq t\leq\ell$ and $Z_{\ell}$ is an orthonormal basis of $\mathrm{span}\{\omega,B\omega,\dotsc,B^{\ell-1}\omega\}$. Then for all $1\leq t\leq N$, there exists a matrix $W_{t}\in\mathbb{R}^{N\times t}$ with orthonormal columns such that
@@
Let $V_{Z}$ and $V_{Z}W_{t}$ be two matrices with orthonormal columns, where their columns form orthonormal bases for the invariant subspaces associated with the $N$ and $t$ smallest eigenvalues of $Z_{\ell}^{\mathsf{T}}BZ_{\ell}$, respectively. By definition, $V=Z_{\ell}V_{Z}$. The gap-independent perturbation result in Proposition 2.4 allows us to conclude (18):
@@
For Eq. 19, the result when $\ell=N$ is straightforward since $\lVert AV\rVert^{2}=\tilde{\lambda}_{N}^{(0)}$. To show the case $\ell>N$, we first note that eigenvalue interlacing implies $\tilde{\lambda}_{N+1}^{(0)}\geq\underline{\sigma}^{2}$. Because of $\|Z_{\ell}^{\mathsf{T}}BZ_{\ell}-Z_{\ell}^{\mathsf{T}}A^{\mathsf{T}}AZ_{\ell}\|\leq\epsilon<\underline{\sigma}^{2}-\tilde{\lambda}_{N}^{(0)}\leq\tilde{\lambda}_{N+1}^{(0)}-\tilde{\lambda}_{N}^{(0)}$, we can apply the second, gap-dependent result in Proposition 2.4 to establish (19).
@@
Let $\overline{\sigma}$ and $\underline{\sigma}$ denote the largest and smallest nonzero singular values of a matrix $A\in\mathbb{R}^{m\times n}$ having null space dimension $N$. Let $V$ be the approximate null space basis returned by Algorithm 1 with $30<\epsilon<\underline{\sigma}^{2}/$ 3 and $N\leq\ell\leq n$. For every $1\leq t\leq N$, there exists a matrix $W_{t}\in\mathbb{R}^{N\times t}$ with orthonormal columns such that
@@
Proof 3.11.
@@
Let $\kappa=\overline{\sigma}/\underline{\sigma}$ be the effective condition number of $A$, and $\epsilon=c\underline{\sigma}^{2}$, where $30<c<1/$ 3. Theorem 3.10 states that, with probability at least $1-3\delta$,
@@
where we used Stirling’s formula and a known bound for Chebyshev polynomials in [28, Section 4.4.1]. Let $BZ_{\ell}=Z_{\ell}T_{\ell}+Q_{\ell+1}E_{\ell+1}$ represent the Lanczos decomposition computed by Algorithm 1. Then, $T_{\ell}-3\epsilon I$ has at least $t$ negative eigenvalues, all originating from Ritz vectors close to the null space. This assertion stems from the fact that the $(N+1)$st smallest eigenvalue of $B$ is at least $\underline{\sigma}^{2}=\epsilon/c>3\epsilon$. Moreover, for $t=N$, the error bound improves to
@@ 4.1 Restarting
Algorithm 1 needs to store the orthonormal basis $Z_{\ell}$ of the Krylov subspace to allow for reorthogonalization (see below) and Ritz vector computations. As the number of iterations $\ell$ required for convergence increases, storing $Z_{\ell}$ becomes challenging due to memory limitations. Additionally, the cost of Rayleigh–Ritz extraction grows significantly with $\ell$. To address these issues, we employ a restart mechanism, akin to the Krylov–Schur method [32, 37].
@@
where $[Z_{\ell}^{(\mathrm{old})},Q_{\ell+1}^{(\mathrm{old})}]$ is a matrix with orthonormal columns, $E_{\ell+1}^{(\mathrm{old})}=[0,\Gamma_{\ell}^{(\mathrm{old})}]$, and
(24) $T_{\ell}^{(\mathrm{old})}=\begin{bmatrix}\Theta^{(\mathrm{old})}&F^{(\mathrm{old})}&&&&\\ (F^{(\mathrm{old})})^{\mathsf{T}}&\Xi_{\ell_{0}+1}^{(\mathrm{old})}&\Gamma_{\ell_{0}+1}^{(\mathrm{old})}&&&\\ &(\Gamma_{\ell_{0}+1}^{(\mathrm{old})})^{\mathsf{T}}&\Xi_{\ell_{0}+2}^{(\mathrm{old})}&\Gamma_{\ell_{0}+2}^{(\mathrm{old})}&&\\ &&\ddots&\ddots&\ddots&\\ &&&(\Gamma_{\ell-2}^{(\mathrm{old})})^{\mathsf{T}}&\Xi_{\ell-1}^{(\mathrm{old})}&\Gamma_{\ell-1}^{(\mathrm{old})}\\ &&&&(\Gamma_{\ell-1}^{(\mathrm{old})})^{\mathsf{T}}&\Xi_{\ell}^{(\mathrm{old})}\\ \end{bmatrix}$
@@
$F^{(\mathrm{new})}=(E_{\ell+1}^{(\mathrm{old})}V_{Z}^{(\mathrm{new})})^{\mathsf{T}}\quad\text{and}\quad\Xi_{\ell_{0}+1}^{(\mathrm{new})}=(Q_{\ell_{0}+1}^{(\mathrm{new})})^{\mathsf{T}}BQ_{\ell_{0}+1}^{(\mathrm{new})},$
@@
$=Z_{\ell_{0}+1}^{(\mathrm{new})}\begin{bmatrix}\Theta^{(\mathrm{new})}&F^{(\mathrm{new})}\\ (F^{(\mathrm{new})})^{\mathsf{T}}&\Xi_{\ell_{0}+1}^{(\mathrm{new})}\end{bmatrix}+Q_{\ell_{0}+2}^{(\mathrm{new})}E_{\ell_{0}+2}^{(\mathrm{new})},$

### v2

@@ 2.3 Principal angles and null space residual
To study the impact of perturbations on subspaces, we recall the definition and basic properties of principal angles. For later purposes, it will be helpful to admit subspaces of unequal dimension, as in [11, Section 6.4.3] and [44].
@@
Note that the principal angles in Definition 2.8 are sorted in increasing order, and it is often the largest principal angle $\angle_{k}(V_{0},V_{1})$ that is of interest. In particular, it is well known [38, Corollary II.4.6] that $\sin\angle_{k}(V_{0},V_{1})=\lVert V_{0}V_{0}^{\mathsf{T}}-V_{1}V_{1}^{\mathsf{T}}\rVert$ holds when $k_{0}=k_{1}$, implying that the orthogonal projectors are close for small $\angle_{k}(V_{0},V_{1})$. The other principal angles give partial information; knowing that some of them are small allows us to conclude that $\mathcal{V}_{0}$ and $\mathcal{V}_{1}$ nearly share a common subspace. In the context of null spaces, the following result establishes a residual bound in terms of principal angles.
@@
Let $W_{t}$ contain the right singular vectors corresponding to the smallest $t$ singular values of $(I-V_{0}V_{0}^{\mathsf{T}})V_{1}$. By [35, Chap. 1, Thm. 4.37], we know that
$\sin\angle_{t}(V_{0},V_{1})=\lVert(I-V_{0}V_{0}^{\mathsf{T}})V_{1}W_{t}\rVert.$
@@
We conclude this section with a useful characterization of the tangents of principal angles from [44, Theorem 3.1].
@@
Let $V_{0},V_{\perp}]$ be an orthonormal matrix, where $V_{0}\in\mathbb{R}^{n\times k_{0}}$. Suppose $\Omega_{1}\in\mathbb{R}^{n\times k_{1}}$ has column full rank. If $\mathrm{rank}(V_{0}^{\mathsf{T}}\Omega_{1})=\mathrm{rank}(\Omega_{1})\leq k_{0}$, then
@@ 3.1 Convergence of single-vector Krylov subspaces for small eigenvalues
In the following, we analyze how well the Krylov subspace generated in line 3 of Algorithm 1 contains approximations to (portions of) the invariant subspace associated with the $N$ smallest eigenvalues of $B$. The analysis mimics analogous results obtained in [22] for low-rank approximation. Compared to classical results [31, Chapter VI], our analysis differs by considering the convergence to a whole subspace and leveraging the randomness of the initial vector. In the following theorem, the main result of this section; $\mathsf{Cheb}_{k}(\lambda)$ denotes the Chebyshev polynomial of degree $k$ defined as
@@
For a symmetric matrix $B\in\mathbb{R}^{n\times n}$ with eigenvalues $\lambda_{1}<\lambda_{2}<\dotsb<\lambda_{N}<\lambda_{N+1}\leq\dotsb\leq\lambda_{n}$, let $V_{N}\in\mathbb{R}^{n\times N}$ be an orthonormal basis of the invariant subspace associated with $\lambda_{1},\dotsc,\lambda_{N}$. Choose integers $t,\ell$ such that $1\leq t\leq N$ and $t\leq\ell\leq n$. Let $K_{\ell}=[\omega,B\omega,\dotsc,B^{\ell-1}\omega]$, where $\omega\in\mathbb{R}^{n}$ is a Gaussian random vector. Then, for any $0<\delta<1$, the inequality
@@
Following [22, Theorem 3], the ability of a (single-vector) Krylov subspace to capture a subspace is closely related to properties of Vandermonde matrices.
@@
where $\mathcal{P}_{t-1}$ denotes the vector space containing all polynomials of degree at most t1t-1. Choosing the Lagrange basis
@@
Because $V_{\perp}^{\mathsf{T}}\omega$ and $V_{N}^{\mathsf{T}}\omega$ are independent Gaussian random vectors, we can use concentration and anti-concentration results for chi-squared random variables from [17, Lemma 1] and [22, Lemma 4], respectively, to conclude that
@@
Our proof crucially relies on a simple but important observation from [22, (1.3)], which establishes the following equivalence between the single-vector Krylov subspace $\mathcal{K}_{\ell}(B,\omega)$ and a block Krylov subspace generated from a specific starting matrix:
(16) $\mathcal{K}_{\ell}(B,\omega)=\mathop{\mathrm{range}}[\omega,B\omega,\dotsc,B^{\ell-1}\omega]=\mathop{\mathrm{range}}[K_{t},BK_{t},\dotsc,B^{\ell-t}K_{t}],$
@@
and an analogous formula holds when $V_{N}$ is replaced by $V_{\perp}$. By assumption, $p(\Lambda_{N})$ is invertible and thus $V_{N}^{\mathsf{T}}KH_{K}$ has rank $t$. Additionally, using that $\mathop{\mathrm{range}}(KH_{K})\subset\mathop{\mathrm{range}}(K)$ and the expression for $\tan\angle_{t}$ from Proposition 2.11, we obtain that
@@
$=\max_{x\in\mathbb{R}^{t}}\frac{\lVert V_{\perp}^{\mathsf{T}}KH_{K}x\rVert}{\lVert V_{N}^{\mathsf{T}}KH_{K}x\rVert}=\max_{x\in\mathbb{R}^{t}}\frac{\lVert p(\Lambda_{\perp})V_{\perp}^{\mathsf{T}}\Omega x\rVert}{\lVert p(\Lambda_{N})V_{N}^{\mathsf{T}}\Omega x\rVert}$
$\leq\frac{\max\limits_{N+1\leq j\leq n}\lvert p(\lambda_{j})\rvert}{\min\limits_{1\leq i\leq N}\lvert p(\lambda_{i})\rvert}\max_{x\in\mathbb{R}^{t}}\frac{\lVert V_{\perp}^{\mathsf{T}}\Omega x\rVert}{\lVert V_{N}^{\mathsf{T}}\Omega x\rVert}=\frac{\max\limits_{N+1\leq j\leq n}\lvert p(\lambda_{j})\rvert}{\min\limits_{1\leq i\leq N}\lvert p(\lambda_{i})\rvert}\lVert V_{\perp}^{\mathsf{T}}\Omega(V_{N}^{\mathsf{T}}\Omega)^{\dagger}\rVert$
$=\frac{\max\limits_{N+1\leq j\leq n}\lvert p(\lambda_{j})\rvert}{\min\limits_{1\leq i\leq N}\lvert p(\lambda_{i})\rvert}\tan\angle_{t}(V_{N},\Omega).$
The proof is completed by setting $p(\lambda)=\mathsf{Cheb}_{\ell-t}\bigl(1+\frac{2(\lambda_{N+1}-\lambda)}{\lambda_{n}-\lambda_{N+1}}\bigr)$ and using that a Chebyshev polynomial is bounded by one in magnitude inside [1,1][-1,1] and monotone outside that interval.
3.2 Convergence of Algorithm 1 with a single initial vector
@@
Under the setting of Theorem 2.6, let $V$ be the approximate null space basis returned after $\ell\geq N$ iterations of Algorithm 1. Let $\tilde{\lambda}_{t}^{(0)}$ be the $t$th smallest eigenvalue of $Z_{\ell}^{\mathsf{T}}A^{\mathsf{T}}AZ_{\ell}$, where $1\leq t\leq\ell$ and $Z_{\ell}$ is an orthonormal basis of $\mathcal{K}_{\ell}(B,\omega)$. Then for all $1\leq t\leq N$, there exists a matrix $W_{t}\in\mathbb{R}^{N\times t}$ with orthonormal columns such that
@@
Let $V_{Z}$ and $V_{Z}W_{t}$ be two matrices with orthonormal columns, where their columns form orthonormal bases for the invariant subspaces associated with the $N$ and $t$ smallest eigenvalues of $Z_{\ell}^{\mathsf{T}}BZ_{\ell}$, respectively. By definition, $V=Z_{\ell}V_{Z}$. The gap-independent perturbation result in Lemma 2.4 allows us to conclude (18):
@@
For Eq. 19, the result when $\ell=N$ is straightforward since $\lVert AV\rVert^{2}=\tilde{\lambda}_{N}^{(0)}$. To show the case $\ell>N$, we first note that eigenvalue interlacing implies $\tilde{\lambda}_{N+1}^{(0)}\geq\underline{\sigma}^{2}$. Because of $\|Z_{\ell}^{\mathsf{T}}BZ_{\ell}-Z_{\ell}^{\mathsf{T}}A^{\mathsf{T}}AZ_{\ell}\|\leq\epsilon<\underline{\sigma}^{2}-\tilde{\lambda}_{N}^{(0)}\leq\tilde{\lambda}_{N+1}^{(0)}-\tilde{\lambda}_{N}^{(0)}$, we can apply the second, gap-dependent result in Lemma 2.4 to establish (19).
@@
Let $\overline{\sigma}$ and $\underline{\sigma}$ denote the largest and smallest nonzero singular values of a matrix $A\in\mathbb{R}^{m\times n}$ having null space dimension $N$. Let $V$ be the approximate null space basis returned by Algorithm 1 with $0<\epsilon<\underline{\sigma}^{2}/3$ and $N\leq\ell\leq n$. For every $1\leq t\leq N$, there exists a matrix $W_{t}\in\mathbb{R}^{N\times t}$ with orthonormal columns such that
@@
The convergence results in Eqs. 20 and 22 consist of two components. The first component (terms in parentheses) captures the convergence of the Krylov subspace method to produce approximations to eigenvalues and eigenvectors of $B$. The second component, discussed in Lemma 3.8, quantifies the quasi-optimality of the Rayleigh–Ritz procedure when applied to $B$ instead of $A^{\mathsf{T}}A$. When measuring the convergence for all zero singular values in Eq. 22, the magnitudes of these two components are both of order $\epsilon^{2}$, and this result is sharp up to a constant for some matrices.
Proof 3.11 (Proof of Theorem 3.10).
@@
Let $\kappa=\overline{\sigma}/\underline{\sigma}$ be the effective condition number of $A$, and $\epsilon=c\underline{\sigma}^{2}$, where $0<c<1/3$. Theorem 3.10 states that, with probability at least $1-3\delta$,
@@
where we used Stirling’s formula and a known bound for Chebyshev polynomials in [31, Section 4.4.1]. Let $BZ_{\ell}=Z_{\ell}T_{\ell}+Q_{\ell+1}E_{\ell+1}$ represent the Lanczos decomposition computed by Algorithm 1. Then, $T_{\ell}-3\epsilon I$ has at least $t$ negative eigenvalues, all originating from Ritz vectors close to the null space. This assertion stems from the fact that the (N+1)(N+1)st smallest eigenvalue of $B$ is at least $\underline{\sigma}^{2}=\epsilon/c>3\epsilon$. Moreover, for $t=N$, the error bound improves to
@@ 4.1 Restarting
Algorithm 1 needs to store the orthonormal basis $Z_{\ell}$ of the Krylov subspace to allow for reorthogonalization (see below) and Ritz vector computations. As the number of iterations $\ell$ required for convergence increases, storing $Z_{\ell}$ becomes challenging due to memory limitations. Additionally, the cost of Rayleigh–Ritz extraction grows significantly with $\ell$. To address these issues, we employ a restart mechanism, akin to the Krylov–Schur method [37, 43, 39].
@@
where $Z_{\ell}^{(\mathrm{old})},Q_{\ell+1}^{(\mathrm{old})}]$ is a matrix with orthonormal columns, $E_{\ell+1}^{(\mathrm{old})}=[0,\Gamma_{\ell}^{(\mathrm{old})}]$, and
(24) $T_{\ell}^{(\mathrm{old})}=\begin{bmatrix}\Theta^{(\mathrm{old})}&(F^{(\mathrm{old})})^{\mathsf{T}}&&&&\\ F^{(\mathrm{old})}&\Xi_{\ell_{0}+1}^{(\mathrm{old})}&(\Gamma_{\ell_{0}+1}^{(\mathrm{old})})^{\mathsf{T}}&&&\\ &\Gamma_{\ell_{0}+1}^{(\mathrm{old})}&\Xi_{\ell_{0}+2}^{(\mathrm{old})}&(\Gamma_{\ell_{0}+2}^{(\mathrm{old})})^{\mathsf{T}}&&\\ &&\ddots&\ddots&\ddots&\\ &&&\Gamma_{\ell-2}^{(\mathrm{old})}&\Xi_{\ell-1}^{(\mathrm{old})}&(\Gamma_{\ell-1}^{(\mathrm{old})})^{\mathsf{T}}\\ &&&&\Gamma_{\ell-1}^{(\mathrm{old})}&\Xi_{\ell}^{(\mathrm{old})}\\ \end{bmatrix}$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F^{(\mathrm{new})}=E_{\ell+1}^{(\mathrm{old})}V_{Z}^{(\mathrm{new})}}\quad\text{and}\quad\Xi_{\ell_{0}+1}^{(\mathrm{new})}=(Q_{\ell_{0}+1}^{(\mathrm{new})})^{\mathsf{T}}BQ_{\ell_{0}+1}^{(\mathrm{new})},$
@@
$=Z_{\ell_{0}+1}^{(\mathrm{new})}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\begin{bmatrix}\Theta^{(\mathrm{new})}&(F^{(\mathrm{new})})^{\mathsf{T}}\\ F^{(\mathrm{new})}&\Xi_{\ell_{0}+1}^{(\mathrm{new})}\end{bmatrix}}+Q_{\ell_{0}+2}^{(\mathrm{new})}E_{\ell_{0}+2}^{(\mathrm{new})},$